\begin{thebibliography}{99}
%\cite{Kachru:2003aw}
\bibitem{Kachru:2003aw}
  S.~Kachru, R.~Kallosh, A.~D.~Linde and S.~P.~Trivedi,
  ``De Sitter vacua in string theory,''
  Phys.\ Rev.\ D {\bf 68} (2003) 046005
 % doi:10.1103/PhysRevD.68.046005
  [hep-th/0301240].
  %%CITATION = doi:10.1103/PhysRevD.68.046005;%%
  %2548 citations counted in INSPIRE as of 19 Aug 2018


 
%\cite{Obied:2018sgi}
\bibitem{Obied:2018sgi}
  G.~Obied, H.~Ooguri, L.~Spodyneiko and C.~Vafa,
  ``De Sitter Space and the Swampland,''
  arXiv:1806.08362 [hep-th].
  %%CITATION = ARXIV:1806.08362;%%
  %31 citations counted in INSPIRE as of 30 Aug 2018
  
  

%\cite{Buscher:1987sk}
\bibitem{Buscher:1987sk}
  T.~H.~Buscher,
  ``A Symmetry of the String Background Field Equations,''
  Phys.\ Lett.\ B {\bf 194} (1987) 59.
 % doi:10.1016/0370-2693(87)90769-6
  %%CITATION = doi:10.1016/0370-2693(87)90769-6;%%
  %746 citations counted in INSPIRE as of 29 Aug 2018


%\cite{Buscher:1987qj}
\bibitem{Buscher:1987qj}
  T.~H.~Buscher,
  ``Path Integral Derivation of Quantum Duality in Nonlinear Sigma Models,''
  Phys.\ Lett.\ B {\bf 201} (1988) 466.
  %doi:10.1016/0370-2693(88)90602-8
  %%CITATION = doi:10.1016/0370-2693(88)90602-8;%%
  %793 citations counted in INSPIRE as of 29 Aug 2018
  
  
  
  
  
  

%\cite{Siegel:1993xq}
\bibitem{Siegel:1993xq}
  W.~Siegel,
  ``Two vierbein formalism for string inspired axionic gravity,''
  Phys.\ Rev.\ D {\bf 47} (1993) 5453
  %doi:10.1103/PhysRevD.47.5453
  [hep-th/9302036].
  %%CITATION = doi:10.1103/PhysRevD.47.5453;%%
  %293 citations counted in INSPIRE as of 29 Aug 2018

%\cite{Siegel:1993th}
\bibitem{Siegel:1993th}
  W.~Siegel,
  ``Superspace duality in low-energy superstrings,''
  Phys.\ Rev.\ D {\bf 48} (1993) 2826
  %doi:10.1103/PhysRevD.48.2826
  [hep-th/9305073].
  %%CITATION = doi:10.1103/PhysRevD.48.2826;%%
  %354 citations counted in INSPIRE as of 29 Aug 2018
  
  
  %\cite{Hull:2009mi}
\bibitem{Hull:2009mi}
  C.~Hull and B.~Zwiebach,
  ``Double Field Theory,''
  JHEP {\bf 0909} (2009) 099
 % doi:10.1088/1126-6708/2009/09/099
  [arXiv:0904.4664 [hep-th]].
  %%CITATION = doi:10.1088/1126-6708/2009/09/099;%%
  %418 citations counted in INSPIRE as of 29 Aug 2018


  %\cite{Hull:2009zb}
\bibitem{Hull:2009zb}
  C.~Hull and B.~Zwiebach,
  %``The Gauge algebra of double field theory and Courant brackets,''
  JHEP {\bf 0909} (2009) 090
 % doi:10.1088/1126-6708/2009/09/090
  [arXiv:0908.1792 [hep-th]].
  %%CITATION = doi:10.1088/1126-6708/2009/09/090;%%
  %243 citations counted in INSPIRE as of 29 Aug 2018
  
  
  
  
  %\cite{Hohm:2010jy}
\bibitem{Hohm:2010jy}
  O.~Hohm, C.~Hull and B.~Zwiebach,
  ``Background independent action for double field theory,''
  JHEP {\bf 1007} (2010) 016
  %doi:10.1007/JHEP07(2010)016
  [arXiv:1003.5027 [hep-th]].
  %%CITATION = doi:10.1007/JHEP07(2010)016;%%
  %296 citations counted in INSPIRE as of 29 Aug 2018
  
  



  
%\cite{Angus:2018mep}
\bibitem{Angus:2018mep}
All the equations of motion of the  closed-string massless sector  coupled to  matter in an $\ODD$-covariant manner can be unified into a single expression, $G_{AB}=8\pi GT_{AB}$:  \\
S.~Angus, K.~Cho and J.~H.~Park,
  ``Einstein Double Field Equations,''
  Eur.\ Phys.\ J.\ C {\bf 78} (2018) no.6,  500
  %doi:10.1140/epjc/s10052-018-5982-y
  [arXiv:1804.00964 [hep-th]].
  %%CITATION = doi:10.1140/epjc/s10052-018-5982-y;%%
  %2 citations counted in INSPIRE as of 29 Aug 2018
  
  
\bibitem{LR}
Equations (\ref{defV1})  and (\ref{defV2})  imply respectively     that    $\big(V_{A}{}^{p},\brV_{A}{}^{\brp}\big)$ is the right and  left inverse of $\big(V^{B}{}_{q},\brV^{B}{}_{\brq}\big)^{T}$. Since  left inverse  is the same as   right inverse in general,  (\ref{defV1})  and (\ref{defV2})  are equivalent.



%\cite{Jeon:2010rw}
\bibitem{Jeon:2010rw}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Differential geometry with a projection: Application to double field theory,''
  JHEP {\bf 1104} (2011) 014 
  %doi:10.1007/JHEP04(2011)014
  [arXiv:1011.1324 [hep-th]].
  %%CITATION = doi:10.1007/JHEP04(2011)014;%%
  %101 citations counted in INSPIRE as of 18 Jun 2016


%\cite{Jeon:2011cn}
\bibitem{Jeon:2011cn}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Stringy differential geometry, beyond Riemann,''
  Phys.\ Rev.\ D {\bf 84} (2011) 044022
  % doi:10.1103/PhysRevD.84.044022
  [arXiv:1105.6294 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.84.044022;%%
  %115 citations counted in INSPIRE as of 23 Mar 2017


%\cite{Morand:2017fnv}
\bibitem{Morand:2017fnv}
  K.~Morand and J.~H.~Park,
  ``Classification of non-Riemannian doubled-yet-gauged spacetime,''
  Eur.\ Phys.\ J.\ C {\bf 77} (2017) no.10,  685
  %doi:10.1140/epjc/s10052-017-5257-z
  [arXiv:1707.03713 [hep-th]].
  %%CITATION = doi:10.1140/epjc/s10052-017-5257-z;%%
  %6 citations counted in INSPIRE as of 21 Aug 2018
 




%\cite{Jeon:2011vx}
\bibitem{Jeon:2011vx}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Incorporation of fermions into double field theory,''
  JHEP {\bf 1111} (2011) 025 
  [arXiv:1109.2035 [hep-th]].
  %%CITATION = ARXIV:1109.2035;%%
  %39 citations counted in INSPIRE as of 20 Feb 2015
%\cite{Jeon:2011sq}
\bibitem{Jeon:2011sq}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Supersymmetric Double Field Theory: Stringy Reformulation of Supergravity,''
  Phys.\ Rev.\ D {\bf 85} (2012) 081501
   [Erratum-ibid.\ D {\bf 86} (2012) 089903]
  [arXiv:1112.0069 [hep-th]].
  %%CITATION = ARXIV:1112.0069;%%
  %49 citations counted in INSPIRE as of 20 Feb 2015

%\cite{Jeon:2012hp}
\bibitem{Jeon:2012hp}
  I.~Jeon, K.~Lee, J.~H.~Park and Y.~Suh,
  ``Stringy Unification of Type IIA and IIB Supergravities under $N=2$ $D=10$ Supersymmetric Double Field Theory,''
  Phys.\ Lett.\ B {\bf 723} (2013) 245
%  doi:10.1016/j.physletb.2013.05.016
  [arXiv:1210.5078 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2013.05.016;%%
  %69 citations counted in INSPIRE as of 29 Aug 2018
 






%\cite{Hohm:2010pp}
\bibitem{Hohm:2010pp}
  O.~Hohm, C.~Hull and B.~Zwiebach,
  ``Generalized metric formulation of double field theory,''
  JHEP {\bf 1008} (2010) 008
  %doi:10.1007/JHEP08(2010)008
  [arXiv:1006.4823 [hep-th]].
  %%CITATION = doi:10.1007/JHEP08(2010)008;%%
  %320 citations counted in INSPIRE as of 22 Aug 2018


%\cite{Cho:2015lha}
\bibitem{Cho:2015lha}
  W.~Cho, J.~J.~Fernández-Melgarejo, I.~Jeon and J.~H.~Park,
  ``Supersymmetric gauged double field theory: systematic derivation by virtue of twist,''
  JHEP {\bf 1508} (2015) 084
  %doi:10.1007/JHEP08(2015)084
  [arXiv:1505.01301 [hep-th]].
  %%CITATION = doi:10.1007/JHEP08(2015)084;%%
  %18 citations counted in INSPIRE as of 29 Aug 2018



  
  %\cite{Geissbuhler:2011mx}
\bibitem{Geissbuhler:2011mx}
  D.~Geissbuhler,
  ``Double Field Theory and N=4 Gauged Supergravity,''
  JHEP {\bf 1111} (2011) 116  
 % doi:10.1007/JHEP11(2011)116
  [arXiv:1109.4280 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2011)116;%%
  %140 citations counted in INSPIRE as of 22 Aug 2018



  
%\cite{Aldazabal:2011nj}
\bibitem{Aldazabal:2011nj}
  G.~Aldazabal, W.~Baron, D.~Marques and C.~Nunez,
  ``The effective action of Double Field Theory,''
  JHEP {\bf 1111} (2011) 052
   Erratum: [JHEP {\bf 1111} (2011) 109]
  %doi:10.1007/JHEP11(2011)052, 10.1007/JHEP11(2011)109
  [arXiv:1109.0290 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2011)052, 10.1007/JHEP11(2011)109;%%
  %170 citations counted in INSPIRE as of 18 Jul 2018










%\cite{Grana:2012rr}
\bibitem{Grana:2012rr}
  M.~Grana and D.~Marques,
  ``Gauged Double Field Theory,''
  JHEP {\bf 1204} (2012) 020 
 % doi:10.1007/JHEP04(2012)020
  [arXiv:1201.2924 [hep-th]].
  %%CITATION = doi:10.1007/JHEP04(2012)020;%%
  %137 citations counted in INSPIRE as of 22 Aug 2018
  
  
  %\cite{Dibitetto:2012rk}
\bibitem{Dibitetto:2012rk}
  G.~Dibitetto, J.~J.~Fernandez-Melgarejo, D.~Marques and D.~Roest,
  ``Duality orbits of non-geometric fluxes,''
  Fortsch.\ Phys.\  {\bf 60} (2012) 1123
  %doi:10.1002/prop.201200078
  [arXiv:1203.6562 [hep-th]].
  %%CITATION = doi:10.1002/prop.201200078;%%
  %102 citations counted in INSPIRE as of 27 Aug 2018


%\cite{Choi:2015bga}
\bibitem{Choi:2015bga}
  K.~S.~Choi and J.~H.~Park,
  ``Standard Model as a Double Field Theory,''
  Phys.\ Rev.\ Lett.\  {\bf 115} (2015) no.17,  171603
  %doi:10.1103/PhysRevLett.115.171603
  [arXiv:1506.05277 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.115.171603;%%
  %7 citations counted in INSPIRE as of 17 Jul 2018



%\cite{Lee:2013hma}
\bibitem{Lee:2013hma}
  K.~Lee and J.~H.~Park,
  ``Covariant action for a string in doubled yet gauged  spacetime,''
  Nucl.\ Phys.\ B {\bf 880} (2014) 134 
  [arXiv:1307.8377 [hep-th]].
  %%CITATION = ARXIV:1307.8377;%%
  %16 citations counted in INSPIRE as of 20 Feb 2015



%\cite{Ko:2015rha}
\bibitem{Ko:2015rha}
  S.~M.~Ko, C.~Melby-Thompson, R.~Meyer and J.~H.~Park,
  ``Dynamics of Perturbations in Double Field Theory \& Non-Relativistic String Theory,''
  JHEP {\bf 1512} (2015) 144
 % doi:10.1007/JHEP12(2015)144
  [arXiv:1508.01121 [hep-th]].
  %%CITATION = doi:10.1007/JHEP12(2015)144;%%
  %13 citations counted in INSPIRE as of 17 Jul 2018

%\cite{Park:2016sbw} 
\bibitem{Park:2016sbw}
  J.~H.~Park,
  ``Green-Schwarz superstring on doubled-yet-gauged spacetime,''
  JHEP {\bf 1611} (2016) 005
%  doi:10.1007/JHEP11(2016)005
  [arXiv:1609.04265 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2016)005;%%
  %11 citations counted in INSPIRE as of 17 Jul 2018


%\cite{Gomis:2000bd}
\bibitem{Gomis:2000bd}
  J.~Gomis and H.~Ooguri,
  ``Nonrelativistic closed string theory,''
  J.\ Math.\ Phys.\  {\bf 42} (2001) 3127
  doi:10.1063/1.1372697
  [hep-th/0009181].
  %107 citations counted in INSPIRE as of 22 Jan 2019



\bibitem{GO}
As a matter of fact,  the Gomis-Ooguri non-relativistic string theory~\cite{Gomis:2000bd} can be derived from the doubled-yet-gauged  string   action~(\ref{stringA}) with a    non-Riemannian    DFT-metric of the type $(1,1)$~\cite{Ko:2015rha}.

\bibitem{Quetient}
The numerators in the quotient groups of (\ref{coset}) are not  target-spacetime  T-duality group:  they act on the local Lorentz vector indices and correspond to  the stabilizer of the defining  relation of the DFT-vielbeins~(\ref{defV2}). 

%\cite{Maharana:1992my}
\bibitem{Maharana:1992my}
  J.~Maharana and J.~H.~Schwarz,
  ``Noncompact symmetries in string theory,''
  Nucl.\ Phys.\ B {\bf 390} (1993) 3
 % doi:10.1016/0550-3213(93)90387-5
  [hep-th/9207016].
  %%CITATION = doi:10.1016/0550-3213(93)90387-5;%%
  %467 citations counted in INSPIRE as of 25 Aug 2018

%\cite{Peeters:2007wn}
\bibitem{Peeters:2007wn}
  K.~Peeters, 
  ``Introducing Cadabra: A Symbolic computer algebra system for field theory problems,''
%\textit{Cadabra}, 
  hep-th/0701238.
  %%CITATION = HEP-TH/0701238;%%
  %137 citations counted in INSPIRE as of 25 Aug 2018







%\cite{Jeon:2011kp}
\bibitem{Jeon:2011kp}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Double field formulation of Yang-Mills theory,''
  Phys.\ Lett.\ B {\bf 701} (2011) 260
  [arXiv:1102.0419 [hep-th]].
  %%CITATION = ARXIV:1102.0419;%%
  %35 citations counted in INSPIRE as of 20 Feb 2015

%\cite{Hohm:2011ex}
\bibitem{Hohm:2011ex}
  O.~Hohm and S.~K.~Kwak,
  ``Double Field Theory Formulation of Heterotic Strings,''
  JHEP {\bf 1106} (2011) 096
  %doi:10.1007/JHEP06(2011)096
  [arXiv:1103.2136 [hep-th]].
  %%CITATION = doi:10.1007/JHEP06(2011)096;%%
  %105 citations counted in INSPIRE as of 23 Aug 2018





%\cite{Berman:2013cli}
\bibitem{Berman:2013cli}
  D.~S.~Berman and K.~Lee,
  ``Supersymmetry for Gauged Double Field Theory and Generalised Scherk-Schwarz Reductions,''
  Nucl.\ Phys.\ B {\bf 881} (2014) 369
    [arXiv:1305.2747 [hep-th]].
  %%CITATION = ARXIV:1305.2747;%%
  %29 citations counted in INSPIRE as of 22 sept. 2015
  


%\cite{Hohm:2014sxa}
\bibitem{Hohm:2014sxa}
  O.~Hohm, A.~Sen and B.~Zwiebach,
  ``Heterotic Effective Action and Duality Symmetries Revisited,''
  JHEP {\bf 1502} (2015) 079
 % doi:10.1007/JHEP02(2015)079
  [arXiv:1411.5696 [hep-th]].
  %%CITATION = doi:10.1007/JHEP02(2015)079;%%
  %20 citations counted in INSPIRE as of 24 Aug 2018
  
  
  %\cite{Malek:2016vsh}
\bibitem{Malek:2016vsh}
  E.~Malek,
  ``From Exceptional Field Theory to Heterotic Double Field Theory via K3,''
  JHEP {\bf 1703} (2017) 057
 % doi:10.1007/JHEP03(2017)057
  [arXiv:1612.01990 [hep-th]].
  %%CITATION = doi:10.1007/JHEP03(2017)057;%%
  %15 citations counted in INSPIRE as of 11 Sep 2018
  



  


%\cite{Aldazabal:2017jhp}
\bibitem{Aldazabal:2017jhp}
  G.~Aldazabal, E.~Andres, M.~Mayo and V.~Penas,
  ``Double Field Theory description of Heterotic gauge symmetry enhancing-breaking,''
  JHEP {\bf 1710} (2017) 046 
%  doi:10.1007/JHEP10(2017)046
  [arXiv:1708.07148 [hep-th]].
  %%CITATION = doi:10.1007/JHEP10(2017)046;%%
  %3 citations counted in INSPIRE as of 23 Aug 2018


\bibitem{footnoteYM}
With (\ref{Wconstraint}),   we can rewrite the Yang--Mills  kinetic term,  
\mbox{{\resizebox{.81\hsize}{!}{$-\textstyle{\frac{1}{8}}\cH^{AC}\cH^{BD}F_{AB}{}^{\dA}F_{CD\dA}=
\textstyle{\frac{1}{2}}P^{AC}\brP^{BD}\cF_{AB}{}^{\dA}\cF_{CD\dA}$}}, where} {\resizebox{.77\hsize}{!}{$\cF_{AB}{}^{\dC}:=\na_{A}W_{B}{}^{\dC}-\na_{B}W_{A}{}^{\dC}+f_{\dA\dB}{}^{\dC}W_{A}{}^{\dA}W_{B}{}^{\dB}$}}~\cite{Choi:2015bga,Jeon:2011kp}.  Irrespective  of (\ref{Wconstraint}) $P_{A}{}^{C}\brP_{B}{}^{D}\cF_{CD}{}^{\dA}$ is diffeomorphism-covariant, but $F_{AB}{}^{\dA}$ is so  if (\ref{Wconstraint}) holds~\cite{Angus:2018mep}.




%\cite{Hassan:1991mq}
\bibitem{Hassan:1991mq}
  S.~F.~Hassan and A.~Sen,
  ``Twisting classical solutions in heterotic string theory,''
  Nucl.\ Phys.\ B {\bf 375} (1992) 103 
 % doi:10.1016/0550-3213(92)90336-A
  [hep-th/9109038].
  %%CITATION = doi:10.1016/0550-3213(92)90336-A;%%
  %207 citations counted in INSPIRE as of 24 Aug 2018



  

%\cite{Bergshoeff:1981um}
\bibitem{Bergshoeff:1981um}
  E.~Bergshoeff, M.~de Roo, B.~de Wit and P.~van Nieuwenhuizen,
  ``Ten-Dimensional Maxwell-Einstein Supergravity, Its Currents, and the Issue of Its Auxiliary Fields,''
  Nucl.\ Phys.\ B {\bf 195} (1982) 97.
  %doi:10.1016/0550-3213(82)90050-5
  %%CITATION = doi:10.1016/0550-3213(82)90050-5;%%
  %536 citations counted in INSPIRE as of 26 Aug 2018

%\cite{Chapline:1982ww}
\bibitem{Chapline:1982ww}
  G.~F.~Chapline and N.~S.~Manton,
  ``Unification of Yang-Mills Theory and Supergravity in Ten-Dimensions,''
  Phys.\ Lett.\ B {\bf 120} (1983) 105.
  % [Phys.\ Lett.\  {\bf 120B} (1983) 105].
%  doi:10.1016/0370-2693(83)90633-0
  %%CITATION = doi:10.1016/0370-2693(83)90633-0;%%
  %631 citations counted in INSPIRE as of 26 Aug 2018


%\cite{Green:1984sg}
\bibitem{Green:1984sg}
  M.~B.~Green and J.~H.~Schwarz,
  ``Anomaly Cancellation in Supersymmetric D=10 Gauge Theory and Superstring Theory,''
  Phys.\ Lett.\  {\bf 149B} (1984) 117.
  %doi:10.1016/0370-2693(84)91565-X
  %%CITATION = doi:10.1016/0370-2693(84)91565-X;%%
  %2635 citations counted in INSPIRE as of 26 Aug 2018





%\cite{Bedoya:2014pma}
\bibitem{Bedoya:2014pma}
  O.~A.~Bedoya, D.~Marques and C.~Nunez,
  ``Heterotic $\alpha^{\prime}$-corrections in Double Field Theory,''
  JHEP {\bf 1412} (2014) 074 
 % doi:10.1007/JHEP12(2014)074
  [arXiv:1407.0365 [hep-th]].
  %%CITATION = doi:10.1007/JHEP12(2014)074;%%
  %38 citations counted in INSPIRE as of 23 Aug 2018


  
%\cite{Coimbra:2014qaa}
\bibitem{Coimbra:2014qaa}
  A.~Coimbra, R.~Minasian, H.~Triendl and D.~Waldram,
  ``Generalised geometry for string corrections,''
  JHEP {\bf 1411} (2014) 160
 % doi:10.1007/JHEP11(2014)160
  [arXiv:1407.7542 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2014)160;%%
  %50 citations counted in INSPIRE as of 28 Aug 2018
  
  
  
 %\cite{Lee:2015kba}
\bibitem{Lee:2015kba}
  K.~Lee,
  ``Quadratic $\alpha^{\prime}$-corrections to heterotic double field theory,''
  Nucl.\ Phys.\ B {\bf 899} (2015) 594 
 % doi:10.1016/j.nuclphysb.2015.08.013
  [arXiv:1504.00149 [hep-th]].
  %%CITATION = doi:10.1016/j.nuclphysb.2015.08.013;%%
  %9 citations counted in INSPIRE as of 23 Aug 2018




%\cite{Hull:2006va}
\bibitem{Hull:2006va}
  C.~M.~Hull,
  ``Doubled Geometry and T-Folds,''
  JHEP {\bf 0707} (2007) 080
 % doi:10.1088/1126-6708/2007/07/080
  [hep-th/0605149].
  %%CITATION = doi:10.1088/1126-6708/2007/07/080;%%
  %220 citations counted in INSPIRE as of 26 Aug 2018


%\cite{Arvanitakis:2017hwb}
\bibitem{Arvanitakis:2017hwb}
  A.~S.~Arvanitakis and C.~D.~A.~Blair,
  ``Unifying Type-II Strings by Exceptional Groups,''
  Phys.\ Rev.\ Lett.\  {\bf 120} (2018) no.21,  211601
 % doi:10.1103/PhysRevLett.120.211601
  [arXiv:1712.07115 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.120.211601;%%
  %3 citations counted in INSPIRE as of 11 Sep 2018


%\cite{Arvanitakis:2018hfn}
\bibitem{Arvanitakis:2018hfn}
  A.~S.~Arvanitakis and C.~D.~A.~Blair,
  ``The Exceptional Sigma Model,''
  JHEP {\bf 1804} (2018) 064
  %doi:10.1007/JHEP04(2018)064
  [arXiv:1802.00442 [hep-th]].
  %%CITATION = doi:10.1007/JHEP04(2018)064;%%
  %4 citations counted in INSPIRE as of 11 Sep 2018



%\cite{Park:2013mpa}
\bibitem{Park:2013mpa}
  {J.\,H.~Park,}
  ``Comments on double field theory and diffeomorphisms,''
  JHEP {\bf 1306} (2013) 098
  [arXiv:1304.5946 [hep-th]].
  %%CITATION = ARXIV:1304.5946;%%
  %21 citations counted in INSPIRE as of 20 Feb 2015


%\cite{Giveon:1991jj}
\bibitem{Giveon:1991jj}
  A.~Giveon and M.~Rocek,
  ``Generalized duality in curved string backgrounds,''
  Nucl.\ Phys.\ B {\bf 380} (1992) 128    
 % doi:10.1016/0550-3213(92)90518-G
[hep-th/9112070].
  %%CITATION = doi:10.1016/0550-3213(92)90518-G;%%
  %169 citations counted in INSPIRE as of 26 Aug 2018


  

\bibitem{33}
For the reduction of the  type II critical   superstring~\cite{Jeon:2012hp,Park:2016sbw} to Minkowskian  four-dimensions (${\hD=10}\rightarrow {D=4}$), we may   consider      ${\Dpr=6}$,   $(3,3)$-type non-Riemannian internal space, as suggested previously   in  \cite{Morand:2017fnv}.  This will   set  the twofold spin groups to be equally ten-dimensional, and   the corresponding  coset structure to be    $\frac{\mathbf{O}(10,10)}{\mathbf{O}(4,6)\times\mathbf{O}(6,4)}$, which    contrasts with   the asymmetric spin groups in (\ref{coset}). 




%\cite{Berman:2019izh}
\bibitem{Berman:2019izh}
  D.~S.~Berman, C.~D.~A.~Blair and R.~Otsuki,
  ``Non-Riemannian geometry of M-theory,''
  arXiv:1902.01867 [hep-th].
  %%CITATION = ARXIV:1902.01867;%%






%\cite{Hohm:2018ybo}
\bibitem{Hohm:2018ybo}
  O.~Hohm and H.~Samtleben,
  ``Leibniz-Chern-Simons Theory and Phases of Exceptional Field Theory,''
  arXiv:1805.03220 [hep-th].
  %%CITATION = ARXIV:1805.03220;%%
  %8 citations counted in INSPIRE as of 02 Mar 2019



\end{thebibliography}

\documentclass[twocolumn,showpacs,preprintnumbers,amsmath,amssymb]{revtex4-1}

\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}
\usepackage[compact]{titlesec}

\usepackage{hyperref}
\newcommand{\half}{{{\textstyle\frac{1}{2}}}}
\newcommand{\quarter}{{{\textstyle\frac{1}{4}}}}
\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation} }
\newcommand{\beqa}{\begin{eqnarray} }
\newcommand{\eeqa}{\end{eqnarray} }
\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}
\newcommand{\bpm}{\begin{pmatrix}}
\newcommand{\epm}{\end{pmatrix}}

\newcommand{\rmd}{{\rm d}}

\newcommand{\rmD}{{\rm D}}

\newcommand{\Apr}{{A^{\prime}{}}}
\newcommand{\Bpr}{{B^{\prime}{}}}
\newcommand{\Dpr}{{D^{\prime}{}}}

\newcommand\cHpr{\cH^{\prime}}
\newcommand\cJpr{\cJ^{\prime}}

\newcommand\hcJ{{\hat{\cal J}}}

\newcommand{\hODD}{\mathbf{O}(\hD,\hD)}
\newcommand{\ODD}{\mathbf{O}(D,D)}
\newcommand{\inODD}{\mathbf{O}(\Dpr,\Dpr)}

\newcommand\So{S_{\scriptscriptstyle{{(0)}}}}
\newcommand\cGo{\cG_{\scriptscriptstyle{{(0)}}}}
\newcommand\brcGo{\brcG_{\scriptscriptstyle{{(0)}}}}
\newcommand\hcGo{\hat{\cG}_{\scriptscriptstyle{{(0)}}}}
\newcommand\hbrcGo{\hat{\brcG}_{\scriptscriptstyle{{(0)}}}}

\newcommand{\brcG}{\bar{\cG}}

\newcommand\Tr{{\rm Tr}}

\newcommand\cA{{\cal A}}

\newcommand\cF{{\cal F}}
\newcommand\cG{{\cal G}}
\newcommand\cH{{\cal H}}

\newcommand\cJ{{\cal J}}

\newcommand\cL{{\cal L}}

\def\tx{\tilde{x}}
\def\ty{\tilde{y}}

\def\na{\nabla}

\def\breta{\bar{\eta}}

\def\brn{{\bar{n}}}
\def\brp{{\bar{p}}}
\def\brq{{\bar{q}}}
\def\brr{{\bar{r}}}

\def\brPhi{{{\bar{\Phi}}}}

\def\brP{\bar{P}}

\def\brV{\bar{V}}

\def\brcF{\bar{\cF}}

\newcommand\hp{\hat{p}}

\newcommand\hA{\hat{A}}
\newcommand\hB{\hat{B}}

\newcommand\hD{\hat{D}}

\newcommand\hV{{\hat{V}}}
\newcommand\hW{{\hat{W}}}

\newcommand\hcH{{\hat{\cH}}}

\def\dA{\dot{A}}
\def\dB{\dot{B}}
\def\dC{\dot{C}}
\def\dD{\dot{D}}
\def\dE{\dot{E}}

\def\dV{\dot{V}}

\def\dcJ{\dot{\cJ}}

\begin{document}

\title{Kaluza--Klein reduction on a maximally non-Riemannian   space is moduli-free}
\author{Kyoungho Cho}
\email{khcho23@sogang.ac.kr}
\author{Kevin Morand}
\email{morand@sogang.ac.kr}
\author{Jeong-Hyuck Park\,}
\email{park@sogang.ac.kr}
 
\affiliation{Department of Physics, Sogang University, 35 Baekbeom-ro, Mapo-gu, Seoul 04107,  Korea}

\begin{abstract}
\centering\begin{minipage}{\dimexpr\paperwidth-7.465cm}
\noindent  We propose  a novel  Kaluza--Klein  scheme  which assumes the internal space to be maximally    non-Riemannian, meaning  that  no  Riemannian  metric can be defined    for any subspace. Its description is  only possible through  Double Field Theory but  not within   supergravity.  We spell out the corresponding  Scherk--Schwarz twistable   Kaluza--Klein ansatz, and point out   that the internal space   prevents rigidly     any    graviscalar moduli.   Plugging the same ansatz into higher-dimensional pure Double Field Theory and also  to a known  doubled-yet-gauged string action, we recover   heterotic supergravity as well as   heterotic  worldsheet   action. In this way, we show that 1) supergravity  and Yang--Mills theory can be unified into higher-dimensional pure Double Field Theory, free of moduli,  and 2) heterotic string theory may have a higher-dimensional non-Riemannian origin.  

\end{minipage}
\end{abstract}

                             
\maketitle
\section{Introduction}
\vspace{5pt}
\noindent Kaluza--Klein theory attempts  to  unify General Relativity  and electromagnetism into higher-dimensional   pure gravity.  Yet, the (aesthetically unpleasing) cylindrical extra  dimension  brings along  an unwanted  additional massless scalar field,   \textit{i.e.~}radion or graviscalar  modulus, which is not observed in nature:  it would  spoil  the Equivalence Principle by appearing on the  right-hand side of the  geodesic equation.   This moduli stabilization   problem is    essentially  rooted in the fact    that  there is no natural scale in pure gravity which would  fix or stabilize  the radius of the cylinder.   The problem persists in modern  string compactifications, in view of  the arbitrary size and shape of an  internal space (mani/coni/orbifold, compact or not).  Turning on fluxes or non-perturbative  corrections might  promise to solve  the problem,  but     such  scenarios   require generically subtle analyses  regarding both the validity of the effective-field-theory approximation   and calculational control~\cite{Kachru:2003aw}  (\textit{c.f.~}\cite{Obied:2018sgi} and references therein for related   current  controversies).

  In this paper  we propose a novel  Kaluza--Klein  scheme to unify Stringy Gravity and Yang--Mills (including Maxwell), which   postulates a non-Riemannian internal space and  consequently  does not suffer from  any moduli problem.   By Stringy Gravity, we mean the string theory effective action of the  NS-NS (or  purely  bosonic) closed-string massless sector,  conventionally represented by the three fields, $\{g_{\mu\nu},B_{\mu\nu},\phi\}$. They transform into one another under  T-duality and hence form a (reducible) $\ODD$  multiplet~\cite{Buscher:1987sk,Buscher:1987qj}.  Furthermore,  within the framework of    Double Field Theory (DFT) initiated in  \cite{Siegel:1993xq,Siegel:1993th,Hull:2009mi,Hull:2009zb,Hohm:2010jy}, $\ODD$ T-duality becomes     the manifest principal  symmetry and      the effective action itself  is   to be  identified as an integral of a   stringy scalar curvature beyond Riemann. The whole  closed-string massless NS-NS sector may  then be   viewed as stringy graviton fields  consisting  of the DFT-dilaton, $d$, and the DFT-metric, $\cH_{AB}$~\cite{Angus:2018mep}. The latter satisfies   two defining properties: 
\be
\ba{ll}
\cH_{AB}=\cH_{BA}\,,\quad&\quad
\cH_{A}{}^{C}\cH_{B}{}^{D}\cJ_{CD}=\cJ_{AB}\,,
\ea
\label{defH}
\ee
\vspace{-3pt}where $\cJ_{AB}={\resizebox{.125\hsize}{!}{${\tiny\mathbf{\Big(\ba{cc} \bf{0}&\bf{1}\\\bf{1}&\bf{0}\ea\Big)}}$}}$ is the $\ODD$ invariant constant metric which can freely raise and lower the $\ODD$ vector indices, $A,B,\cdots$ (capital letters).   A pair of  symmetric  projection matrices  are then defined,
\be
\ba{ll}
P_{AB}=\half(\cJ_{AB}+\cH_{AB})\,,\quad&
\brP_{AB}=\half(\cJ_{AB}-\cH_{AB})\,,
\ea
\ee
while their  `square roots'  give twofold DFT-vielbeins, 
\be
\ba{ll}
P_{AB}=V_{A}{}^{p}V_{B}{}^{q}\eta_{pq}\,,\quad&\quad
\brP_{AB}=\brV_{A}{}^{\brp}\brV_{B}{}^{\brq}\breta_{\brp\brq}\,,
\ea
\ee
which satisfy their own  defining properties, 
\be
\ba{lll}
V_{Ap}V^{A}{}_{q}=\eta_{pq}\,,\quad&~~
\brV_{A\brp}\brV^{A}{}_{\brq}=\breta_{\brp\brq}\,,\quad&~~
V_{Ap}\brV^{A}{}_{\brq}=0\,,
\ea
\label{defV1}
\ee
or equivalently~\cite{LR}
\be
V_{A}{}^{p}V_{B}{}^{q}\eta_{pq}+\brV_{A}{}^{\brp}\brV_{B}{}^{\brq}\breta_{\brp\brq}=\cJ_{AB}\,.
\label{defV2}
\ee
Here $\eta_{pq}$ and $\breta_{\brp\brq}$ are  local  Lorentz invariant    metrics   characterizing  the twofold spin groups which  have  dimensions $\Tr(P)$ and $\Tr(\brP)$ separately.  They are distinguished by  unbarred and barred small  letter   indices.

The above stringy graviton fields constitute  the    diffeomorphic DFT-Christoffel symbols~\cite{Jeon:2010rw,Jeon:2011cn,Morand:2017fnv},
{\small{\[
\ba{ll}
\!\Gamma_{CAB}=&\!\!{2\left(P\partial_{C}P\brP\right)_{[AB]}
+2\left({{\brP}_{[A}{}^{D}{\brP}_{B]}{}^{E}}-{P_{[A}{}^{D}P_{B]}{}^{E}}\right)\partial_{D}P_{EC}}\\
{}&\!-4\!\left(\textstyle{\frac{P_{C[A}P_{B]}{}^{D}}{P_{F}{}^{F}-1}}+\textstyle{\frac{\brP_{C[A}\brP_{B]}{}^{D}}{\brP_{F}{}^{F}-1}}\right)\!\left(\partial_{D}d+(P\partial^{E}P\brP)_{[ED]}\right),
\ea
\label{Gammao}
\]}}and, consequently  with $\na_{A}:=\partial_{A}+\Gamma_{A}$,    twofold local Lorentz spin connections~\cite{Jeon:2011vx,Jeon:2011sq,Jeon:2012hp},
\be
\ba{l}
\Phi_{Apq}=V^{B}{}_{p}\na_{A}V_{Bq}=V^{B}{}_{p}\!\left(\partial_{A}V_{Bq}+\Gamma_{AB}{}^{C}V_{Cq}\right),\\
\brPhi_{A\brp\brq}=\brV^{B}{}_{\brp}\na_{A}\brV_{B\brq}=\brV^{B}{}_{\brp}\!\left(\partial_{A}\brV_{B\brq}+\Gamma_{AB}{}^{C}\brV_{C\brq}\right).
\ea
\label{PhibrPhi}
\ee
These connections form  covariant curvatures, Ricci and scalar~\cite{Jeon:2011cn,Jeon:2012hp}, of which the latter  can be constructed as
\be
\ba{l}
\So=(P^{AC}P^{BD}-\brP^{AC}\brP^{BD})S_{ABCD}\,,\\
\cGo=\cF_{ABpq}V^{Ap}V^{Bq}+\half\Phi_{Apq}\Phi^{Apq}\,,\\
\brcGo=\brcF_{AB\brp\brq}\brV^{A\brp}\brV^{B\brq}+\half\brPhi_{A\brp\brq}\brPhi^{A\brp\brq}\,,
\ea
\ee
where, with {\resizebox{.8\hsize}{!}{\,${R_{CDAB}=\partial_{A}\Gamma_{BCD}+\Gamma_{AC}{}^{E}\Gamma_{BED}-(A\leftrightarrow B)}$}}, 
\[
\ba{l}
\!S_{ABCD}:=\half\left(R_{ABCD}+R_{CDAB}-\Gamma^{E}{}_{AB}\Gamma_{ECD}\right)\,,\\
\!\cF_{ABpq}:=\na_{A}\Phi_{Bpq}-\na_{B}\Phi_{Apq}+\Phi_{Ap}{}^{r}\Phi_{Brq}-\Phi_{Bp}{}^{r}\Phi_{Arq}\,,\\
\!\brcF_{AB\brp\brq}:=\na_{A}\brPhi_{B\brp\brq}-\na_{B}\brPhi_{A\brp\brq}+\brPhi_{A\brp}{}^{\brr}\brPhi_{B\brr\brq}-\brPhi_{B\brp}{}^{\brr}\brPhi_{A\brr\brq}\,.
\ea
\]
While $\So$ coincides  exactly  with  the    well-known expression of  scalar curvature written in terms of $d, \cH_{AB}$~\cite{Hohm:2010pp}, 
\[
\ba{l}
\!\!\So=\textstyle{\frac{1}{8}}\cH^{AB}\partial_{A}\cH_{CD}\partial_{B}\cH^{CD}+\half\cH^{AB}
\partial^{C}\cH_{AD}\partial^{D}\cH_{BC}\\
\!-\partial_{A}\partial_{B}\cH^{AB}
+4\cH^{AB}(\partial_{A}\partial_{B}d-\partial_{A}d\partial_{B}d)+4\partial_{A}\cH^{AB}\partial_{B}d\,,
\ea
\]
the other two accommodate  the vielbeins and read~\cite{Cho:2015lha}
\be
\ba{rl}
+\cGo&\!\!=\half\So+2\partial_{A}\partial^{A}d-2\partial_{A}d\partial^{A}d+\half\partial_{A}V_{Bp}\partial^{A}V^{Bp}\,,\\\
-\brcGo&\!\!=\half\So-2\partial_{A}\partial^{A}d+2\partial_{A}d\partial^{A}d-\half\partial_{A}\brV_{B\brp}\partial^{A}\brV^{B\brp}\,.
\ea
\label{cGobrcGo}
\ee
Clearly their differences  would vanish upon the section condition, $\partial_{A}\partial^{A}\equiv0$, but they  provide   precisely the known `missing' pieces  in  the Scherk--Schwarz reduction of DFT  while    relaxing the  section condition  on  the internal space, as  previously  added by hand in \cite{Geissbuhler:2011mx,Aldazabal:2011nj,Grana:2012rr,Dibitetto:2012rk}.  Hence, the above vielbein formalism generates these crucial terms in a natural manner. In particular,  $\cGo\!-\brcGo$ matches  the action  adopted in  \cite{Grana:2012rr}. Below we  focus on computing   $+\cGo$ and $-\brcGo$ separately in higher dimensions   with  the Scherk--Schwarz  twisted non-Riemannian  Kaluza--Klein ansatz. \\


\section{\mbox{\!\!\!\!\!Moduli-free non-Riemannian Kaluza--Klein ansatz}}
\vspace{5pt}
\noindent With ${\hD=\Dpr+D}$, the  DFT Kaluza--Klein ansatz~\cite{Morand:2017fnv}  breaks $\hODD$ into $\inODD\times\ODD$ and takes the form:
\be
\ba{ll}
\hcH={\resizebox{.5\hsize}{!}{$\exp\!\left[\hW\right]\left(\ba{cc}\cHpr&0\\0&\cH\ea\right)\exp\!\left[\hW^{T}\right]$}}\,,
\quad&~~\hcJ={\resizebox{.17\hsize}{!}{$\left(\ba{cc}\cJpr&0\\0&\cJ\ea\right)$}}\,.
\ea
\label{ansatz}
\ee
In our notation,     hatted, primed, and unprimed symbols refer to  the    ambient, internal, and external  spaces   respectively. In particular,  the ambient doubled coordinates  split   into the internal and external ones as   
\be
\ba{ll}
z^{\hA}=(y^{A^{\prime}},x^{A})\,,\quad&\quad
\partial_{\hA}=(\partial_{A^{\prime}},\partial_{A})\,.
\ea
\label{split}
\ee
In (\ref{ansatz}),   $\hW$   is an off-block-diagonal $\mathfrak{so}(\hD,\hD)$ element which  satisfies 
\be
\hW\hcJ+\hcJ\hW^{T}=0\,,
\ee
and takes  the form,
\be
\!\hW\!=\!{\resizebox{.95\hsize}{!}{$
\left(\!\ba{cc}0&-W^{{\mathbf{c}}}\\W&\!0\ea\!\right)\!,~ 
W^{{\mathbf{c}}}_{~\Apr}{}^{\!A\!}:=W^{A}{}_{\!\Apr}= \cJ^{AB}\cJpr_{\Apr\Bpr}W_{B}{}^{\Bpr},$}}
\ee
where  the $2D\times 2\Dpr$    block,  $W_{A}{}^{\Apr}$,  should  meet~\cite{Choi:2015bga}  
\be
\ba{ll}
W_{A}{}^{\Apr}W^{A\Bpr}=0\,,\qquad&\quad W^{A}{}_{\Apr}\partial_{A}=0\,.
\ea
\label{Wconstraint} 
\ee
This    condition sets half of the components to be trivial, truncates the exponential, 
\[
\exp\!\left[\hW\right]=1+\hW+\half\hW^{2}\,,
\]
and makes  the above  ansatz    consistent with the ordinary  Kaluza--Klein  ansatz in supergravity.    Moreover, crucially for the purpose of the present work,  the   ansatz~(\ref{ansatz}) can   accommodate   non-Riemannian geometry in which the Riemannian metric cannot be defined,   see \cite{Morand:2017fnv} for classification and \cite{Lee:2013hma,Ko:2015rha,Park:2016sbw} for earlier examples  including    the attainment of  the Gomis-Ooguri non-relativistic string~\cite{Gomis:2000bd,GO}.  

 Henceforth,  we focus on a specific  internal space  of which the DFT-metric is fully $\inODD$ symmetric and   maximally  non-Riemannian, namely $(\Dpr,0)$ type as classified in   \cite{Morand:2017fnv}:
\be
\cHpr_{\Apr\Bpr}\equiv\cJpr_{\Apr\Bpr}\,.
\label{internal}
\ee     
In general,  from the defining relations~(\ref{defH}),  the infinitesimal variation of  any  DFT-metric should satisfy
\be
\delta \cHpr=P^{\prime}\delta\cHpr\brP^{\prime}+\brP^{\prime}\delta\cHpr P^{\prime}\,.
\ee
Meanwhile, the  particular   choice of the internal space~(\ref{internal}) implies ${P^{\prime}=\cJpr}$ and ${\brP^{\prime}=0}$.  Thus,  the   fluctuation must be trivial:   no graviscalar modulus  can be generated and the non-Riemannian internal space is \textit{rigid}, 
\be
\delta \cHpr_{\Apr\!\Bpr}=0\,.
\label{rigid}
\ee  
In fact,   (\ref{internal}) sets the ``twofold" internal spin group to be   $\inODD\times\mathbf{O}(0,0)$,  such that   the coset structures  of the internal and the ambient  DFT-metrics, $\cHpr$, $\hcH$, are `trivial' and `heterotic'  respectively, if $\cH$ is Riemannian~(\ref{Riemann}), 
\be
{\resizebox{.95\hsize}{!}{$
\frac{\inODD}{\inODD\times\mathbf{O}(0,0)}=\mathbf{1}\,,\quad
\frac{\hODD}{\mathbf{O}(\Dpr+1,\hD-1)\times\mathbf{O}(D-1,1)}\,.$}}
\label{coset}
\ee
The latter has  dimension ${D^{2}+2D\Dpr}$, which  matches the total degrees of the  external DFT-metric, $\cH_{AB}$  (\ref{Riemann}),  and the gravivector,  $W_{A}{}^{\Apr}$~(\ref{Wconstraint})~\cite{Quetient}, \textit{c.f.~}\cite{Maharana:1992my}.   

The corresponding DFT-vielbeins are~\cite{Morand:2017fnv}
\be
\ba{ll}
\hV_{\hA\hp}\!=\!\exp\!\left[\hW\right]\!\left(\ba{cc}
V^{\prime}_{A^{\prime}p^{\prime}}&0\\
0&V_{Ap}\ea\right),&
\hat{\brV}_{\hA\hat{\brp}}\!=\!\exp\!\left[\hW\right]\left(\ba{c}
0\\
\brV_{A\brp}\ea\right),
\ea
\label{ansatzV}
\ee
where $V^{\prime}_{A^{\prime}p^{\prime}}$ is now an invertible   `square' matrix,
 \[
V^{\prime}_{A^{\prime}p^{\prime}}V^{\prime}_{B^{\prime}}{}^{p^{\prime}}=\cJ^{\prime}_{A^{\prime}B^{\prime}}\,.
\]

\section{Reduction to heterotic DFT}
\vspace{5pt}
\noindent Before inserting  the ansatz~(\ref{ansatz}), (\ref{ansatzV}) into the ambient scalar curvatures~(\ref{cGobrcGo}), or $\hcGo,\hbrcGo$ (all hatted),  we   perform a Scherk--Schwarz twist over the internal space.  Following \cite{Cho:2015lha}, we  introduce a twisting matrix, {\resizebox{.12\hsize}{!}{${U_{\dA}{}^{A^{\prime}}(y)}$}}, which depends on the internal coordinates only and is an $\inODD$  element satisfying 
\be
U_{\dA}{}^{A^{\prime}}U_{\dB}{}^{B^{\prime}}\cJpr_{A^{\prime}B^{\prime}}=\dcJ_{\dA\dB}\,,
\ee
where $\dcJ_{\dA\dB}$ coincides with $\cJpr_{A^{\prime}B^{\prime}}$ numerically: both are $\inODD$ invariant constant metrics.  Essentially   the twist converts   all the primed indices to  dotted ones, 
\be
\ba{ll}
\! \!\!W_{A}{}^{\Apr\!}(z)=W_{A}{}^{\dA}(x)U_{\dA}{}^{A^{\prime}\!}(y)\,,&\!
V^{\prime A^{\prime}}{}_{\! p^{\prime}}(z)=\dV{}^{\dA}{}_{p^{\prime}}(x)U_{\dA}{}^{A^{\prime}\!}(y)\,,
\ea
\label{twist1}
\ee
where now  $\dV_{\dA p^{\prime}}\dV_{\dB}{}^{p^{\prime}}=\dcJ_{\dA\dB}$.    We further put
\be
\ba{lll}
\hat{d}(z)=\lambda(y)+d(x)\,,\quad&\quad
V_{Ap}(x)\,,\quad&\quad \brV_{A\brp}(x)\,,
\ea
\label{trunc}
\ee
such that the external fields are independent of the internal coordinates:  $\partial_{A^{\prime}}d=0$\,,  $\partial_{A^{\prime}}V_{Bp}=0$\,, $\partial_{A^{\prime}}\brV_{B\brp}=0$\,.

Finally, we impose the standard section condition on the external space, $\partial_{A}\partial^{A}=0$,  and the twistability conditions on the internal space separately~\cite{Grana:2012rr},\cite{Cho:2015lha}:
\be
\ba{ll}
\partial_{\Apr}\lambda-\half U^{\dA}{}_{\Apr}\partial_{\Bpr}U_{\dA}{}^{\Bpr}=0\,,\quad&\quad
f_{[\dA\dB}{}^{\dE}f_{\dC]\dD\dE}=0\,,
\ea
\label{twist2}
\ee
where  $f_{\dA\dB\dC}=3\big(\partial_{\Bpr}U_{[\dA}{}^{\Apr}\big)U_{\dB}{}^{\Bpr}U_{\dC]\Apr}$ which we require  to be  constant.  

After straightforward yet tedious computation  -- assisted by a computer algebra system~\cite{Peeters:2007wn} and through intermediate expressions like (4.14) in   \cite{Cho:2015lha} -- we obtain our main result: with the above  Scherk--Schwarz twisted Kaluza--Klein ansatz substituted,  the higher-dimensional   scalar curvatures~(\ref{cGobrcGo})  reduce precisely to (\textit{c.f.}~\cite{Jeon:2011kp,Hohm:2011ex,Grana:2012rr,Berman:2013cli,Hohm:2014sxa,Malek:2016vsh,Aldazabal:2017jhp})
\be
\ba{lll}
\!\!+2\hcGo&\!=\!&-2\hbrcGo+\textstyle{\frac{1}{3}}f_{\dA\dB\dC}f^{\dA\dB\dC}\,,\\
\!\!-2\hbrcGo&\!=\!&\So-\textstyle{\frac{1}{4}}\cH^{AC}\cH^{BD}
F_{AB}{}^{\dA}F_{CD\dA}\\
{}&{}&-\textstyle{\frac{1}{12}}\cH^{AD}\cH^{BE}\cH^{CF}\omega_{ABC}
\omega_{DEF}\\
{}&{}&+\textstyle{\frac{1}{2}}\cH^{AD}\cH^{BE}\cH^{CF}\omega_{ABC}\cH_{[D}{}^{G}\partial_{E}\cH_{F]G}\,,
\ea
\label{MAIN}
\ee
where we set Yang--Mills  and Chern--Simons terms,
\be
\ba{l}
F_{AB}{}^{\dC}=\partial_{A}W_{B}{}^{\dC}-\partial_{B}W_{A}{}^{\dC}+f_{\dA\dB}{}^{\dC}W_{A}{}^{\dA}W_{B}{}^{\dB}\,,\\
\omega_{ABC}=3W_{[A}{}^{\dA}\partial_{B}W_{C]\dA}+f_{\dA\dB\dC}W_{A}{}^{\dA}W_{B}{}^{\dB}W_{C}{}^{\dC}\,,
\ea
\label{Fomega}
\ee
of which the $\ODD$ indices are totally skew-symmetric,   $F_{AB}{}^{\dC}=F_{[AB]}{}^{\dC}$, $
\omega_{ABC}=\omega_{[ABC]}$, and further from (\ref{Wconstraint}),
\be
\ba{ll}
F^{AB}{}_{\dC}\partial_{A}=0\,,\quad&\quad\omega^{ABC}\partial_{A}=0\,.
\ea
\ee

The higher  $\hD$-dimensional  diffeomorphisms  generated by the  standard DFT  Lie derivative    give rise, with a  twisted  parameter, $\hat{\xi}{}^{\hA}=(\Lambda^{\dA}U_{\dA}{}^{\Apr},\xi^{A})$, to the $D$-dimensional diffeomorphisms plus Yang--Mills gauge symmetry, \textit{c.f.}~\cite{Grana:2012rr,Cho:2015lha}, 
\be
\ba{rll}
\delta W_{A}{}^{\dA}&=&\xi^{C}\partial_{C}W_{A}{}^{\dA}
+\left(\partial_{A}\xi^{B}-\partial^{B}\xi_{A}\right)W_{B}{}^{\dA}\\
{}&{}&+\partial_{A}\Lambda^{\dA}+f^{\dA}{}_{\dB\dC}W_{A}{}^{\dB}\Lambda^{\dC}\,,\\
\delta\cH_{AB}&=&\xi^{C}\partial_{C}\cH_{AB}+2\partial_{[A}\xi_{C]}\cH^{C}{}_{B}+
2\partial_{[B}\xi_{C]}\cH_{A}{}^{C}\\
{}&{}&+\big(W_{[A}{}^{\dA}\partial_{C]}\Lambda_{\dA}\big)\cH^{C}{}_{B}+\big(
W_{[B}{}^{\dA}\partial_{C]}\Lambda_{\dA}\big)\cH_{A}{}^{C}\,,\\
\delta d&=&\xi^{C}\partial_{C}d-\half\partial_{C}\xi^{C}\,.
\ea
\label{SYMMETRY}
\ee
In each transformation above, the first line with $\xi^{A}$ is  the (external) diffeomorphic DFT Lie derivative  and  the second   with $\Lambda^{\dA}$   is the  (internal) Yang--Mills gauge symmetry.  In fact, every single term in (\ref{MAIN}) is (external) diffeomorphism-invariant~\cite{footnoteYM}.

It is worth while  to note the only difference between    $2\hcGo$  and $-2\hbrcGo$:   the former contains  a DFT-cosmological constant term~\cite{Jeon:2011cn},    $\textstyle{\frac{1}{3}}e^{-2d}f_{\dA\dB\dC}f^{\dA\dB\dC}$, 
but the latter  does not.   As  anticipated in \cite{Hassan:1991mq,Hohm:2014sxa},  our result~(\ref{MAIN}) is manifestly symmetric for $\ODD$  as well as  any  subgroup of $\inODD$ which stabilizes the structure constant, $f_{\dA\dB\dC}$.

Finally, if we adopt the well-known Riemannian parametrization of the DFT-metric,  
\be
\ba{ll}
\cH_{AB}=\left(\ba{cc}g^{-1}&-g^{-1}B\\
Bg^{-1}&g-Bg^{-1}B\ea\right),\quad&\quad e^{-2d}=\sqrt{-g}e^{-2\phi}\,,
\ea
\label{Riemann}
\ee
and solve both the external section condition, $\partial_{A}\partial^{A}=0$, and   (\ref{Wconstraint}), by letting    $\partial_{A}=(\tilde{\partial}^{\mu},\partial_{\nu})\equiv(0,\partial_{\nu})$ and $W_{A}{}^{\dA}\equiv(0,W_{\nu}{}^{\dA})$,  our main result~(\ref{MAIN})  reproduces    the heterotic supergravity  action~\cite{Bergshoeff:1981um,Chapline:1982ww}, up to total derivatives, 
\be
\ba{l}
\displaystyle{
\int -2e^{-2d}\brcGo}\\
\displaystyle{\!\!
=\!\!\int\!\!\!\sqrt{-g}e^{-2\phi}\big(R+\!4\partial_{\mu}\phi\partial^{\mu}\phi-\!\textstyle{\frac{1}{12}}\tilde{H}_{\lambda\mu\nu}\tilde{H}^{\lambda\mu\nu\!}-\!\quarter F_{\mu\nu\dC}F^{\mu\nu\dC}
\big)},
\ea
\ee
where 
\be
\tilde{H}_{\lambda\mu\nu}=3\partial_{[\lambda}B_{\mu\nu]}-\omega_{\lambda\mu\nu}\,,
\ee 
which is invariant under the Yang--Mills gauge symmetry~(\ref{SYMMETRY}), or  specifically in components,
\[
\ba{ll}
\delta W_{\mu}{}^{\dA}=\partial_{\mu}\Lambda^{\dA}+f^{\dA}{}_{\dB\dC}W_{\mu}{}^{\dB}\Lambda^{\dC}\,,\quad&\quad
\delta B_{\mu\nu}=W_{[\mu}{}^{\dA}\partial_{\nu]}\Lambda_{\dA}\,.
\ea
\] 
However, we have not pinned down the gauge group to be either $\mathbf{E}_{8}\times \mathbf{E}_{8}$ or $\mathbf{SO}(32)$.  
Further, as the starting scalar curvatures~(\ref{cGobrcGo}) are at most two-derivative,  our final  result~(\ref{MAIN})  lacks the  higher-derivative gravitational Chern--Simons term~\cite{Green:1984sg}, \textit{c.f.~}\cite{Bedoya:2014pma,Coimbra:2014qaa,Lee:2015kba}.\\

\section{Reduction to heterotic string}
\vspace{5pt}
\noindent  Henceforth we discuss the worldsheet aspect of the DFT background~(\ref{ansatz}) involving   the maximally  non-Riemannian internal space. The  analysis of  the string  propagating on the most  general  non-Riemannian background was carried out  in \cite{Morand:2017fnv}, with the conclusion  that string becomes generically  (anti-)chiral.    To review this and  apply it to the current   non-Riemannian background~(\ref{ansatz}),  we recall the  doubled-yet-gauged string action~\cite{Hull:2006va},\,\cite{Lee:2013hma} (\textit{c.f.~}\cite{Park:2016sbw,Arvanitakis:2017hwb,Arvanitakis:2018hfn}),
\be
S_{\scriptstyle{\rm{string}}}={\textstyle{\frac{1}{4\pi\alpha^{\prime}}}}{\displaystyle{\int}}\rmd^{2}\sigma~\cL_{\scriptstyle{\rm{string}}}\,,
\label{stringA}
\ee
which contains a generic DFT-metric, 
\[
\cL_{\scriptstyle{\rm{string}}}=-\half\sqrt{-h}h^{\alpha\beta}\rmD_{\alpha}z^{\hA}\rmD_{\beta}z^{\hB}\hcH_{\hA\hB}
-\epsilon^{\alpha\beta}\rmD_{\alpha}z^{\hA}\cA_{\beta\hA}\,.
\label{stringaction}
\]
With  ambient doubled coordinates,  $z^{\hA}=(\tilde{z}_{\hat{\mu}},z^{\hat{\nu}})$, and an auxiliary gauge potential, $\cA_{\alpha}{}^{\hA}$,  a covariant derivative is introduced, 
\be
\ba{ll}
{\rmD_{\alpha} z^{\hA}:=\partial_{\alpha}z^{\hA}-\cA_{\alpha}{}^{\hA}}\,,\qquad&\quad
\cA_{\alpha}{}^{\hA}\partial_{\hA}=0\,.
\ea
\label{rmDz}
\ee
While this action is completely covariant with respect to the desired symmetries like      $\ODD$ T-duality, Weyl symmetry, and  worldsheet as well as  target-spacetime diffeomorphisms, it also     realizes  concretely   the assertion  that  doubled coordinates in DFT are actually gauged and each gauge orbit in the doubled coordinate system corresponds to a single physical point~\cite{Park:2013mpa}.  Specifically,  with the choice of the section, $\partial_{\hA}=(\tilde{\partial}^{\hat{\mu}},\partial_{\hat{\nu}})\equiv(0,\partial_{\hat{\nu}})$,  the condition on the auxiliary gauge potential~(\ref{rmDz})  is  solved by $\cA_{\alpha}{}^{\hA}=(\cA_{\alpha\hat{\mu}},0)$. Therefore,    half of the doubled coordinates, namely  the  tilde coordinates, are gauged,
\[
{\rmD_{\alpha} z^{\hA}=\big(\partial_{\alpha}\tilde{z}_{\hat{\mu}}-\cA_{\alpha\hat{\mu}}\,,\,\partial_{\alpha} z^{\hat{\nu}}\big)}\,.
\]
Upon the Riemannian background~(\ref{Riemann}),  the auxiliary gauge potential  appears quadratically in  the action~(\ref{stringA}). Then after integrating it out,   one  recovers   the standard undoubled string action with   $g_{\mu\nu}$ and $B_{\mu\nu}$.  On the  other hand, upon a generic   non-Riemannian DFT background which is characterized by two non-negative integers, $(n,\brn)$, the  components of the  auxiliary field,  $\cA_{\alpha}{}^{\hA}$,  appear linearly for  $n$ and $\brn$ directions, playing    the role of Lagrange multipliers. Consequently,   string becomes  chiral and anti-chiral over the $n$ and $\brn$ directions respectively~\cite{Morand:2017fnv}.

Now, for  the present   non-Riemannian background~(\ref{ansatz}),  in accordance with the decomposition of  the  ambient space~(\ref{split}),  we put  $y^{\Apr}=(\ty_{\mu^{\prime}},y^{\nu^{\prime}})$  for the internal doubled coordinates,   $x^{A}=(\tx_{\mu},x^{\nu})$    for the external ones,  and   $W_{\mu}{}^{\Apr}=(W_{\mu\nu^{\prime}},\tilde{W}_{\mu}{}^{\nu^{\prime}})$  for the gravivector. Further,   for the ambient  gauge potential,    $\cA_{\alpha}{}^{\hA}=(\cA_{\alpha}{}^{\Apr}, \cA_{\alpha}{}^{A})$, we set  the internal and the external ones as   $\cA_{\alpha}{}^{\Apr}=(A_{\alpha\mu^{\prime}},0)$ and $\cA_{\alpha}{}^{A}=(A_{\alpha\mu},0)$.  With this preparation and  (\ref{ansatz}),  (\ref{Riemann}),  the string action~(\ref{stringA}) becomes quadratic in the combination of the  gauge potential components,  $A_{\alpha\mu}-\tilde{W}_{\mu}{}^{\mu^{\prime}}A_{\alpha\mu^{\prime}}$,  and linear in $A_{\alpha\mu^{\prime}}$.  Hence,   the former leads to a Gaussian integral and  the latter plays the role of a Lagrange multiplier.  After all,   the   doubled-yet-gauged string action~(\ref{stringA})  reduces to 
$\frac{1}{2\pi\alpha^{\prime}}
\displaystyle{\int}\rmd^{2}\sigma~\cL_{\scriptstyle{\rm{Het}}}$,  with   
\be
\ba{lll}
\!\cL_{\scriptstyle{\rm{Het}}}\!&\!\!=\!\!\!&
-\half\sqrt{-h}h^{\alpha\beta}\partial_{\alpha}x^{\mu}\partial_{\beta}x^{\nu}
g_{\mu\nu}
+\half\epsilon^{\alpha\beta}\partial_{\alpha}x^{\mu}
\partial_{\beta}x^{\nu}B_{\mu\nu}\\
{}&{}&+\half\epsilon^{\alpha\beta}
(\partial_{\alpha}\ty_{\mu^{\prime}}+\partial_{\alpha}x^{\mu}W_{\mu\mu^{\prime}})
(\partial_{\beta}y^{\mu^{\prime}}+\partial_{\beta}x^{\lambda}\tilde{W}_{\lambda}{}^{\mu^{\prime}})\\
{}&{}&+\half\epsilon^{\alpha\beta}
(\partial_{\alpha}\tx_{\mu}-W_{\mu\Apr}\partial_{\alpha}y^{\Apr})\partial_{\beta}x^{\mu}\,.
\ea
\label{LHet}
\ee
Furthermore, $\partial_{\alpha}y^{\mu^{\prime}}+\partial_{\alpha}x^{\mu}\tilde{W}_{\mu}{}^{\mu^{\prime}}$ must  be \textit{chiral}:
\be
\left(h^{\alpha\beta}+\textstyle{\frac{1}{\sqrt{-h}}}\epsilon^{\alpha\beta}\right)\!\left(\partial_{\beta}y^{\mu^{\prime}}+\partial_{\beta}x^{\mu}\tilde{W}_{\mu}{}^{\mu^{\prime}}\right)=0\,.
\label{chiral}
\ee
Especially when the gravivector, $W_{\mu}{}^{\dA}$,  is trivial,   the second and  third lines in (\ref{LHet})  merge into  a   known  topological term, $\half\epsilon^{\alpha\beta}\partial_{\alpha}\tilde{z}_{\hat{\mu}}
\partial_{\beta}z^{\hat{\mu}}$~\cite{Giveon:1991jj,Hull:2006va},  while  (\ref{chiral})  gets  simplified to make    the internal coordinates  chiral: 
\be
y^{\mu^{\prime}}(\tau,\sigma)=y^{\mu^{\prime}}(0,\tau+\sigma)\,.
\ee
This implies -- at least classically~\cite{Morand:2017fnv,Ko:2015rha} --  that a closed string subject to a periodic boundary condition cannot vibrate  in the internal space.  In summary, upon the DFT background~(\ref{ansatz}) with the maximally non-Riemannian internal space and trivial gravivector,  the external  fields, $x^{\mu}$,  are described by the usual string action corresponding to the first line in (\ref{LHet}), and  consist  of the equal number of  left-moving and   right-moving sectors. On the other hand,      the internal  fields, $y^{\mu^{\prime}}$, become all chiral, which   agrees with the rigidity~(\ref{rigid}).\\

\section{Conclusion}
\vspace{5pt}
\noindent The maximally non-Riemannian DFT background specified by  the DFT-metric,  $\cH^{\prime}_{\Apr\Bpr}=\cJ^{\prime}_{\Apr\Bpr}$~(\ref{internal}),    is   singled out  to be  completely $\inODD$ symmetric and rigid: it does not admit any  linear fluctuation, $\delta\cH^{\prime}_{\Apr\Bpr}=0$~(\ref{rigid}), nor graviscalar moduli, and the coset structure is trivial~(\ref{coset}).  It is the most symmetric vacuum in DFT.  \\
\indent For  the  DFT Kaluza--Klein ansatz, (\ref{ansatz}) and (\ref{ansatzV}), we  set the internal space to be maximally non-Riemannian, performed   a  Scherk--Schwarz twist~(\ref{twist1}),~(\ref{twist2}), and  computed the ambient higher $(D^{\prime}+D)$-dimensional DFT scalar curvatures, $+\hcGo$ and  $-\hbrcGo$~(\ref{cGobrcGo}), which lead to the   $\ODD$-manifest  formulation of the  non-Abelian heterotic supergravity~(\ref{MAIN}).  Only the former, $+\hcGo$,  contains a DFT-cosmological constant.  In this way,  \textit{supergravity and Yang-Mills theory can be unified into a higher-dimensional pure Stringy Gravity, free of  moduli.}  \\   
\indent Plugging   the \textit{same} non-Riemannian Kaluza--Klein ansatz~(\ref{ansatz}) into  the doubled-yet-gauged string action~(\ref{stringA})  may reproduce  the usual  heterotic  string action,  
\[{\resizebox{.97\hsize}{!}{\,${\half\int_{\Sigma}}
\big({-\sqrt{-h}h^{\alpha\beta}}
g_{\mu\nu}+\epsilon^{\alpha\beta}B_{\mu\nu}\big)\partial_{\alpha}x^{\mu}
\partial_{\beta}x^{\nu}+\epsilon^{\alpha\beta}\partial_{\alpha}\tilde{z}_{\hat{\mu}}
\partial_{\beta}z^{\hat{\mu}}\,,$}}
\] 
with chiral internal coordinates,  $\big(h^{\alpha\beta}+\frac{\epsilon^{\alpha\beta}}{\sqrt{-h}}\big)\partial_{\beta}y^{\mu^{\prime}}=0$.   This is indeed  the case when the Yang--Mills sector is trivial.  Therefore, \textit{heterotic string theory may have  a higher-dimensional origin with non-Riemannian internal space.}

Our analyses have been focused on the bosonic sectors, both on the target-spacetime and   on the world-sheet.  Inclusion of fermions is  highly  desired, \textit{c.f.~}\cite{Jeon:2011sq,Jeon:2012hp,Park:2016sbw}.   
We also leave the exploration of  the worldsheet aspect of the somewhat mysterious  `relaxed'  section condition designed  for the  Scherk--Schwarz twist~\cite{Geissbuhler:2011mx,Aldazabal:2011nj,Grana:2012rr,Dibitetto:2012rk,Cho:2015lha} for future work with  nontrivial gravivector, $W_{A}{}^{\Apr}\neq0$,  for the action~(\ref{LHet}).  
Uplift of the Standard Model of particle physics  coupled to DFT~\cite{Choi:2015bga} to higher dimensions (bottom-up) would be also of interest, as well as applications to  string compactifications  (top-down), possibly on other types of non-Riemannian internal  spaces~\cite{33}.\\


\noindent\textit{Note added.}  \\
Motivated by the \href{https://arxiv.org/abs/1808.10605}{first version of this work on arXiv},  Berman,  Blair, and Otsuki  explored  non-Riemannian geometries in M-theory~\cite{Berman:2019izh}. In particular, they pointed out that the maximally non-Riemannian $E_{8}$ background naturally  realizes the topological  (hence `moduli-free') phase  of  Exceptional Field Theory~\cite{Hohm:2018ybo}. \\

\noindent\textit{Acknowledgments.}\\  {We  wish to thank Stephen Angus,  Kanghoon Lee for useful comments,  and     Wonyoung Cho for  valuable help at the early stage of the project.  This work  was  supported by  the National Research Foundation of Korea   through  the Grants,  NRF-2016R1D1A1B01015196  and NRF-2018H1D3A1A01030137~(Korea Research Fellowship Program).}
\hfill

\end{document}